# ContrastScore: Towards Higher Quality, Less Biased, More Efficient Evaluation Metrics with Contrastive Evaluation

### v1

@@ Abstract
Evaluating the quality of generated text automatically remains a significant challenge. Conventional reference-based metrics have been shown to exhibit relatively weak correlation with human evaluations. Recent research advocates the use of large language models (LLMs) as source-based metrics for natural language generation (NLG) assessment. While promising, LLM-based metrics, particularly those using smaller models, still fall short in aligning with human judgments. In this work, we introduce ContrastScore, a contrastive evaluation metric designed to enable higher-quality, less biased, and more efficient assessment of generated text. We evaluate ContrastScore on two NLG tasks: machine translation and summarization. Experimental results show that ContrastScore consistently achieves stronger correlation with human judgments than both single-model and ensemble-based baselines. Notably, ContrastScore based on Qwen 3B and 0.5B even outperforms Qwen 7B, despite having only half as many parameters, demonstrating its efficiency. Furthermore, it effectively mitigates common evaluation biases such as length and likelihood preferences, resulting in more robust automatic evaluation.
@@ 1 Introduction
Evaluating the quality of automatically generated text remains a fundamental challenge in natural language processing (NLP) and, in some cases, is nearly as difficult as generating the text itself. Traditional evaluation methods primarily rely on reference-based metrics such as BLEU, ROUGE, and METEOR  Papineni et al. (2002); Lin (2004); Banerjee and Lavie (2005); Shen et al. (2022); Pan et al. (2024); Kalyan (2024), which are inadequate as they compare generated text to human-written references based on surface-level lexical overlap, often failing to capture semantic adequacy and fluency. Embedding-based metrics like BERTScore Zhang et al. (2019) improve upon lexical approaches by leveraging contextualized representations for better semantic alignment. However, they remain sensitive to domain shifts, depend on high-quality references, and exhibit relatively weak correlations with human judgments Zhao et al. (2023).
To address these limitations, recent research has shifted towards source-based evaluation methods, leveraging large language models (LLMs). Metrics such as BARTScore (Yuan et al., 2021) and GPTScore (Fu et al., 2024) assess generation quality using a single model’s probability distribution, while prompt-based LLM evaluators process structured prompts to evaluate text based on predefined criteria Que et al. (2024) or conduct comparative judgments without an explicit rubric Liusie et al. (2023). Although these methods show promise, their effectiveness is inherently constrained by the underlying model’s capacity. Larger models generally perform better but are computationally expensive and incur high API costs (Larionov and Eger, 2024). Conversely, smaller models often have reduced capacity, resulting in unreliable evaluations. Furthermore, existing LLM-based metrics are susceptible to biases (Deutsch et al., 2022; Sun et al., 2022; Liu et al., 2024), including length bias and over-reliance on likelihood, which may not always align with human evaluation criteria. These challenges highlight the need for novel evaluation methods that are both efficient and reliable, balancing model capacity, accessibility, and evaluation robustness (Chen and Eger, 2023; Zhao et al., 2024).
@@
Inspired by contrastive principles, we introduce ContrastScore, an evaluation metric that leverages structured disagreement between two models of differing capacities. Unlike conventional LLM-based evaluation methods that rely solely on a single model’s likelihood estimates, ContrastScore incorporates a weaker auxiliary model as a contrastive signal, dynamically adjusting probability scores to improve alignment with human judgements. Specifically, ContrastScore is built on a discrepancy-based probability formulation that measures the absolute difference between the probabilities assigned by two models. By utilizing discrepancies between a stronger (expert) model and a weaker (amateur) model, it generates more calibrated and robust evaluation scores.
We conduct extensive experiments on two widely employed text generation tasks, namely, machine translation and summarization, to evaluate the effectiveness of ContrastScore. Our evaluation compares ContrastScore against established metrics and various baseline models, including single-model and ensemble-based approaches. Experimental results show that ContrastScore achieves a higher correlation with human judgments, outperforming both single-model and ensemble-based methods. Notably, ContrastScore using smaller models (Qwen 3B, Qwen 0.5B) even surpasses Qwen7B with 5.2% of improvement in summarization despite having only half the parameters, demonstrating its efficiency. Furthermore, ContrastScore based on smaller models substantially enhances evaluation speed, providing at least a 1.5-fold increase in processing speed compared to a single larger model across both the Qwen and Llama families. Additionally, it effectively mitigates length and likelihood biases, enhancing robustness in automatic evaluation. By incorporating contrastive principles into evaluation, ContrastScore paves the way for a new paradigm of more robust and efficient text evaluation.
To summarize, the contributions of our paper are four-fold:
  • We propose a simple yet highly effective additive-based formulation for contrastive evaluation, leveraging structured model discrepancies to produce more calibrated and reliable evaluation scores.
@@
  • Contrastive evaluation strongly correlates with human judgments, outperforming existing source-based metrics while effectively addressing likelihood and length biases, which are prevalent in automatic evaluation.
  • ContrastScore achieves significantly faster evaluation compared to larger single-model approaches, improving inference speed by at least 1.5 times while using only half the parameters.
@@ 2 Related Work
LLM-Based Evaluation.
With the rise of LLMs, evaluation has shifted towards using LLMs as direct evaluators of text generation quality. Among LLM-based approaches, BARTScore [^1] The original BARTScore used BART as a language model, but this can be replaced by any more recent LLM such as GPT4 (Fu et al., 2024). Yuan et al. (2021) has emerged as a widely used and reliable metric due to its flexibility and strong correlation with human judgments. It leverages a pretrained model to compute log-likelihood scores for generated text, making it a strong indicator of text quality.
Beyond BARTScore, another prominent evaluation paradigm involves using LLMs as evaluators via prompting. This approach, commonly referred to as LLM-as-a-Judge, directly queries LLMs with carefully designed prompts to assess the quality of generated text. One notable framework following this paradigm is G-Eval Liu et al. (2023), which systematically prompts LLMs to evaluate text across multiple quality attributes (e.g. coherence and consistency), allowing for a more structured and adaptable assessment. Similarly, ChatGPT-based evaluation Gao et al. (2023) explores the use of ChatGPT as a human-like summarization evaluator, demonstrating its ability to approximate human judgments. These methods leverage the implicit knowledge and reasoning abilities of LLMs to provide fine-grained evaluations, making them well-suited for tasks where traditional metrics fall short. ChatEval Chan et al. (2024) introduces a multi-LLM evaluation framework based on a multi-agent debate system. Unlike single-model evaluators, ChatEval orchestrates multiple LLM-based agents, each assuming different roles and perspectives.
Shortcoming of LLM-based Evaluation.
Prompting-based evaluation techniques suffer from prompt sensitivity, where small changes in the input prompt can lead to significantly different results. This sensitivity raises concerns about result consistency, as even slight rewordings or rephrasings of the prompt can alter the model’s response Leiter and Eger (2024). Furthermore, these approaches rely on the capabilities of the specific LLM used, meaning that while they may perform well when evaluated with state-of-the-art models, the reliability of their assessments deteriorates when used with weaker models.
More broadly, LLM-based evaluation methods inherit the biases and limitations of the underlying model. One key issue is likelihood bias, where LLMs tend to favor sentences with higher likelihood scores based on their internal language modeling prior, regardless of the task’s requirements or the actual quality of the generated text Ohi et al. (2024). This bias leads to inflated scores for fluent but incorrect or hallucinated outputs, undermining the evaluation’s reliability. Furthermore, LLMs may exhibit biases inherited from their training data, leading to systematic errors in assessing factuality, coherence, or stylistic appropriateness.
Contrastive Decoding.
@@
Contrastive methods have recently gained attention in text generation, where the differences between models of varying capacities are leveraged to improve output quality Su et al. (2022). Rather than relying on ensemble methods or majority voting, contrastive decoding explicitly uses a weaker model as a contrastive signal to refine or steer the output of a stronger model. This approach has demonstrated success in mitigating common generation issues such as repetition, overconfidence, and incoherence. Li et al. (2023) utilizes a large ‘expert’ model and a smaller ‘amateur’ model. Instead of directly following the expert’s probabilities, contrastive decoding selects tokens that maximize the log-probability difference between the two models. This results in text that is more fluent, diverse, and factually accurate. Moreover, contrastive principles have been extended to model architecture optimization. Shi et al. (2024) propose the Self-Contrast Mixture-of-Experts (SCMoE) framework, where different expert pathways within the same model act as internal contrastive filters, which improves reasoning capabilities and overall model performance, eliminating the need for external weaker models. Note that contrastive decoding is also related to language model arithmetic in which arithmetic combinations of LLMs are considered for adjusted generation (Dekoninck et al., 2024). This includes subtraction, which leads to output favored by one LLM but disfavored by the other, which could be leveraged in adversarial settings (Zhang and Eger, 2024).
@@
Inspired by these advancements, as described in §3.3, we extend contrastive principles beyond generation to the domain of evaluation. Instead of refining model outputs, we propose ContrastScore—a novel metric with a different formulation from prior contrastive decoding work. By leveraging a discrepancy-based probability framework that quantifies the absolute difference between the probabilities assigned by two models, our approach generates contrastive signals that enable more calibrated and reliable evaluation scores.
@@ 3 Methodology
The primary goal of this study is to introduce a contrastive evaluation framework, ContrastScore. Current generative metrics for text evaluation are predominantly based on a single language model. However, their performance is highly dependent on the capabilities of the underlying model, leading to performance disparities, where metrics based on smaller or medium-sized models typically underperform those using larger models. Additionally, these single-model approaches are inherently affected by the biases of the underlying model such as likelihood bias, where the model may over-prefer fluent but less accurate outputs, or pretrained biases that can influence evaluation outcomes in unintended ways.
To address this, we introduce ContrastScore, a method that leverages the structured disagreement between two models of different capacities to produce a bias-aware, dynamically calibrated probability score for evaluation. Rather than relying solely on the probability of a single model, ContrastScore incorporates an auxiliary model as a contrastive signal, allowing for a more refined probability adjustment that improves alignment with human judgments.
@@ 3.1 Generative Evaluation
$\text{Score}(\textbf{h}\mid d,\mathcal{S})=\sum_{t=1}^{m}\log P(h_{t}\mid\textbf{h}_{<t},\mathcal{S},\theta),$ (1)
where $\mathcal{S}$ represents the supplementary text, which may consist of the source text s (in a source-based setting) or the reference output r (in a reference-based setting). $\theta$ denotes the model parameters, and $d$ is the description of the task instruction.
@@ 3.2 Contrastive Decoding
[…]
Here, $V_{\text{head}}(h_{<t})$ selects the top most likely tokens $x_{t}$ under the expert model given history $h_{<t}$. During next token prediction, CD-Score effectively only considers the most likely tokens of the expert model as candidates, and then assigns modified probabilities $\log{\frac{p_{\text{EXP}}^{t}}{p_{\text{AMA}}^{t}}}$ to them. The intuition behind CD-Score is to consider the most likely expert tokens as continuation for generation but then choose those tokens among them which the amateur might not favor (e.g., repetitive tokens).
Unfortunately, CD-Score is unsuitable for evaluation because every generated sequence h containing a token $h_{t}$ not in $V_{\text{head}}(h_{<t})$ would receive a score of $\infty$- ∞, essentially being ruled out as a candidate, even though it should have been considered as part of the sequence under evaluation. As an alternative, one might consider the division score $\log{\frac{p_{\text{EXP}}^{t}}{p_{\text{AMA}}^{t}}}$ itself as an objective for evaluation. However, as pointed out by Li et al. (2023), this approach suffers from penalizing many standard text sequences as unlikely, for which both expert and amateur would hold the same probabilities. For example, if both models assign a probability of 0.9 or 0.1 to a token, the score remains the same, making it unable to distinguish between confidently correct and incorrect predictions. As a result, this approach fails to highlight important distinctions between tokens that should matter in evaluation. The division score also becomes instable especially when the amateur probability is close to zero, distorting the evaluation score.
@@ 3.3 ContrastScore
To address this issue for evaluation, we propose a subtraction-based contrastive formulation with a scaling factor $\gamma\in[0,1]$ that downweights the amateur model. In this way, we naturally retain all tokens in the input sequence for evaluation, as well as a property that favors generations preferred by the expert model, similar to CD-Score. Importantly, the subtraction-based formulation ensures that the scores of individual tokens remain comparable across positions—unlike the division-based approach, where small variations in the amateur model’s likelihood can lead to disproportionately large and unstable scores. By using a small $\gamma$, we control the influence of the amateur model’s probability and mitigate such instability, yielding more reliable evaluation signals.
Empirically, we choose a formula for contrastive evaluation that leverages the absolute value distance between the expert and the scaled down amateur:
$\text{ContrastScore}=\sum_{t=1}^{m}\log\left(\left|p_{\text{EXP}}^{t}-\gamma p_{\text{AMA}}^{t}\right|\right)$ (5)
At each time step $t$, ContrastScore rewards tokens $h_{t}$ for which expert and (downweighted) amateur generation probabilities are maximally distinct (as the absolute value is a distance function between two distributions). This would ensure that ContrastScore leverages the strengths of the expert model but removes the limitations of the amateur model; in practice, it might for example lean toward generations that are less repetitive, which the amateur model tends to favor. Note, however, that there are probability ranges where the scaled down amateur would be preferred over the expert — for example, when $p^{t}_{\text{EXP}}=0$, then our formula would still assign the scaled probability of the amateur to $h_{t}$. With $\gamma=1$, ContrastScore would be indifferent between the expert and the amateur but indiscriminately favor disagreement between them, an undesirable behavior similar to the division score above.

### v2

@@ Abstract
Recent advances in automatic evaluation of natural language generation have increasingly relied on large language models as general-purpose metrics. While effective, these approaches often require high-capacity models, which introduce substantial computational costs, and remain susceptible to known evaluation pathologies, such as over-reliance on likelihood. We introduce ContrastScore, a contrastive evaluation paradigm that builds on the widely used BARTScore formulation by comparing token-level probabilities between a stronger and a weaker model. Instead of relying on single-model likelihoods or prompt-based judgments, ContrastScore captures disagreement between models to better reflect confidence and uncertainty in generation quality. Empirical results on summarization and machine translation benchmarks show that ContrastScore, instantiated with paired moderate-scale models across both Qwen and LLaMA families, consistently outperforms larger alternatives, such as Qwen 7B and LLaMA 8B, in correlation with human ratings. In addition to improving evaluation quality, ContrastScore significantly reduces susceptibility to likelihood bias, offering a more robust and cost-effective alternative to larger LLM-based evaluation methods.  [^1] Source code is available at https://github.com/sandywangxiao/ContrastScore.
ContrastScore: Towards Higher Quality, Less Biased, More Efficient Evaluation Metrics with Contrastive Evaluation
Xiao Wang1, Daniil Larionov2, Siwei Wu1, Yiqi Liu1, Steffen Eger3, Nafise Sadat Moosavi4, Chenghua Lin1 1The University of Manchester, 2University of Mannheim, 3University of Technology Nuremberg, 4The University of Sheffield {xiao.wang-26,siwei.wu-2,yiqi.liu-6}@postgrad.manchester.ac.uk daniil.larionov@uni-mannheim.de, steffen.eger@utn.de n.s.moosavi@sheffield.ac.uk, chenghua.lin@manchester.ac.uk
@@ 1 Introduction
Evaluating the quality of automatically generated text remains a fundamental challenge in natural language processing (NLP) and, in some cases, is nearly as difficult as generating the text itself. Traditional evaluation methods primarily rely on reference-based metrics such as BLEU, ROUGE, and METEOR  Papineni et al. (2002); Lin (2004); Banerjee and Lavie (2005); Shen et al. (2023); Pan et al. (2024); Kalyan (2024), which are inadequate as they compare generated text to human-written references based on surface-level lexical overlap, often failing to capture semantic adequacy and fluency. Embedding-based metrics like BERTScore Zhang et al. (2020) improve upon lexical approaches by leveraging contextualized representations for better semantic alignment. However, they remain sensitive to domain shifts, depend on high-quality references, and exhibit relatively weak correlations with human judgments Zhao et al. (2023). To address these limitations, recent research has shifted towards source-based evaluation methods, leveraging large language models (LLMs). Metrics such as BARTScore (Yuan et al., 2021) and GPTScore (Fu et al., 2024) assess generation quality using a single model’s probability distribution, while prompt-based LLM evaluators process structured prompts to evaluate text based on predefined criteria Que et al. (2024) or conduct comparative judgments without an explicit rubric Liusie et al. (2024). Although these methods show promise, their effectiveness is inherently constrained by the underlying model’s capacity. Larger models generally perform better but are computationally expensive and incur high API costs (Larionov and Eger, 2024). Conversely, smaller models often have reduced capacity, resulting in unreliable evaluations. Furthermore, existing LLM-based metrics are susceptible to biases (Deutsch et al., 2022; Sun et al., 2022; Lu et al., 2023; Liu et al., 2024; Hong et al., 2025), including over-reliance on likelihood, which may not align with human evaluation criteria. These challenges highlight the need for novel evaluation methods that are both efficient and reliable, balancing model capacity, accessibility, and evaluation robustness (Chen and Eger, 2023; Zhao et al., 2024).
@@
Inspired by contrastive principles, we introduce ContrastScore, an evaluation metric that leverages structured disagreement between two models of differing capacities. Unlike conventional LLM-based evaluation methods that rely solely on a single model’s likelihood estimates, ContrastScore incorporates a weaker auxiliary model as a contrastive signal, dynamically adjusting probability scores to improve alignment with human judgements. Specifically, ContrastScore is built on a discrepancy-based probability formulation that measures the absolute difference between the probabilities assigned by two models. By utilizing discrepancies between a stronger (expert) model and a weaker (amateur) model, it generates more calibrated and robust evaluation scores. We conduct extensive experiments on two widely employed text generation tasks, namely, machine translation and summarization, to evaluate the effectiveness of ContrastScore. Our evaluation compares ContrastScore against established metrics and various baseline models, including single-model and ensemble-based approaches. Experimental results show that ContrastScore achieves a higher correlation with human judgments, outperforming both single-model and ensemble-based methods. Notably, ContrastScore using smaller models (Qwen 3B, Qwen 0.5B) even surpasses Qwen7B with 7.0% of improvement in summarization despite having only half the parameters, demonstrating its efficiency. Furthermore, ContrastScore based on smaller models substantially enhances evaluation speed, providing at least a 1.5-fold increase in processing speed compared to a single larger model across both the Qwen and LLaMA families. Additionally, it effectively mitigates the likelihood biase, enhancing robustness in automatic evaluation. By incorporating contrastive principles into evaluation, ContrastScore paves the way for a new paradigm of more robust and efficient text evaluation.
The contributions of our paper are four-fold:
  • We propose a simple yet highly effective difference-based formulation for contrastive evaluation, leveraging structured model discrepancies to produce more calibrated and reliable evaluation scores.
@@
  • Contrastive evaluation strongly correlates with human judgments, outperforming single-model and ensemble methods while effectively addressing likelihood bias, which are prevalent in automatic evaluation.
  • ContrastScore achieves significantly faster evaluation compared to larger single-model approaches, improving inference speed by at least 1.5 times while using only half the parameters, while maintaining comparable or even superior performance.
@@ 2 Related Work
Automatic Evaluation of Text Generation.
Automatic evaluation metrics for text generation can broadly be categorized into task-specific and general-purpose approaches. Early task-specific metrics such as BLEU Papineni et al. (2002), ROUGE Lin (2004), and CHRF Popović (2015) rely on surface-level n-gram overlap with a reference. While originally designed for tasks like machine translation and summarization, these metrics often fail to reflect actual semantic similarity due to the variability of valid natural language expressions. To overcome these limitations, embedding-based metrics, such as BERTScore Zhang et al. (2020) and MoverScore Zhao et al. (2019), were introduced to compute token-level or document-level similarity using contextualized embeddings . Further improvements came from task-specific metrics fine-tuned on human ratings, such as COMET Rei et al. (2020), COMET-KIWI Rei et al. (2022), BLEURT Sellam et al. (2020), and Prism Thompson and Post (2020). These models learn to directly predict human preferences but often generalize poorly across domains and tasks. General-purpose evaluation has increasingly shifted toward leveraging LLMs as evaluators. Prompt-based approaches such as G-Eval Liu et al. (2023), ChatEval Chan et al. (2024); Hong et al. (2025), and GPT-based judge systems treat LLMs as reference-free annotators that directly assess text quality via carefully crafted prompts. While these methods often achieve high correlation with human judgments, they typically rely on proprietary models like GPT-4, making them costly, and sensitive to prompt phrasing Leiter and Eger (2024). An alternative line of work explores probability-based evaluation metrics, including BARTScore Yuan et al. (2021) and GPTScore Fu et al. (2024). BARTScore and GPTScore share a common formulation: they compute the conditional log-likelihood of a generation given context, using a pretrained language model as the scoring function. However, these evaluation methods inherit the biases and limitations of the underlying model Ohi et al. (2024).
Contrastive Paradigm.
@@
Contrastive methods have recently gained attention in text generation, where the differences between models of varying capacities are leveraged to improve output quality Su et al. (2022); O’Brien and Lewis (2023); Li et al. (2023). Rather than relying on ensemble methods or majority voting, contrastive decoding explicitly uses a weaker model as a contrastive signal to refine or steer the output of a stronger model. This approach has demonstrated success in mitigating common generation issues such as repetition, overconfidence, and incoherence. Moreover, contrastive principles have been extended to model architecture optimization. Shi et al. (2024) propose the Self-Contrast Mixture-of-Experts (SCMoE) framework, where different expert pathways within the same model act as internal contrastive filters, which improves reasoning capabilities and overall model performance. Note that contrastive decoding is also related to language model arithmetic in which arithmetic combinations of LLMs are considered for adjusted generation (Dekoninck et al., 2024). This includes subtraction, which leads to output favored by one LLM but disfavored by the other, which could be leveraged in adversarial settings (Zhang and Eger, 2024).
@@
Inspired by these advancements, as described in §3.3, we extend contrastive principles beyond generation to the domain of evaluation. Instead of refining model outputs, we propose ContrastScore—a novel metric with a different formulation from prior contrastive decoding work.
@@ 3 Methodology
The primary goal of this study is to overcome limitations of single-model evaluators, such as capacity bias and over-reliance on model-specific likelihoods, by capturing where strong and weak models disagree in token-level probabilities. This section first reviews generative likelihood-based evaluation and contrastive decoding, then presents the design and formulation of ContrastScore.
@@ 3.1 Generative Evaluation
$\text{Score}(\textbf{h}|d,\mathcal{S})=\sum_{t=1}^{m}w_{t}\log P(h_{t}|\textbf{h}_{<t},\mathcal{S},\theta)$ (1)
where $\mathcal{S}$ represents the supplementary text, which may consist of the source text s (in a source-based setting) or the reference output r (in a reference-based setting). $d$ denotes the evaluation dimension (e.g., fluency), $w_{t}$  [^2] Following prior studies Yuan et al. (2021); Qin et al. (2023), $w_{t}$ is weighed equally for all tokens in this work. refers to the token-level weight, and $\theta$ is the model parameters.
@@ 3.2 Contrastive Decoding
$\text{CD-Score}=\begin{cases}\log{\frac{p_{\text{EXP}}^{t}}{p_{\text{AMA}}^{t}}}&\text{if }h_{t}\in V_{\text{head}}(h_{<t})\\ -\infty&\text{else}\end{cases}$ (4)
Here, $V_{\text{head}}(h_{<t})$ selects the top most likely tokens $x_{t}$ under the expert model given history $h_{<t}$. During next token prediction, CD-Score effectively only considers the most likely tokens of the expert model as candidates, and then assigns modified logits $\log{\frac{p_{\text{EXP}}^{t}}{p_{\text{AMA}}^{t}}}$ to them. The intuition behind CD-Score is to consider the most likely expert tokens as continuation for generation but then choose those tokens among them which the amateur might not favor (e.g., repetitive tokens).
Unfortunately, CD-Score is unsuitable for evaluation because every generated sequence h containing a token $h_{t}$ not in $V_{\text{head}}(h_{<t})$ would receive a score of -$\infty$, essentially being ruled out as a candidate, even though it should have been considered as part of the sequence under evaluation. As an alternative, one might consider the division score $\log{\frac{p_{\text{EXP}}^{t}}{p_{\text{AMA}}^{t}}}$ itself as an objective for evaluation. However, as pointed out by Li et al. (2023), this approach suffers from penalizing many standard text sequences as unlikely, for which both expert and amateur would hold the same probabilities. For example, if both models assign a probability of 0.9 or 0.1 to a token, the score remains the same, making it unable to distinguish between confidently correct and incorrect predictions. As a result, this approach fails to highlight important distinctions between tokens that should matter in evaluation. The division score also becomes instable especially when the amateur probability is close to zero.
@@ 3.3 ContrastScore
To address this issue for evaluation, we propose a subtraction-based contrastive formulation with a scaling factor $\gamma\in[0,1]$ that downweights the amateur model. In this way, we naturally retain all tokens in the input sequence for evaluation, as well as a property that favors generations preferred by the expert model, similar to CD-Score. Importantly, the subtraction-based formulation ensures that the scores of individual tokens remain comparable across positions—unlike the division-based approach, where small variations in the amateur model’s likelihood can lead to disproportionately large and unstable scores. By using a small $\gamma$, we control the influence of the amateur model’s probability and mitigate such instability, yielding more reliable evaluation signals. Empirically, we choose a formula for contrastive evaluation that leverages the absolute value distance between the expert and the scaled down amateur:
$\text{ContrastScore}=\sum_{t=1}^{m}w_{t}\log\left(\left|p_{\text{EXP}}^{t}-\gamma p_{\text{AMA}}^{t}\right|\right)$ (5)
At each time step $t$, ContrastScore rewards tokens for which expert and (downweighted) amateur generation probabilities are maximally distinct (as the absolute value is a distance function between two distributions). This would ensure that ContrastScore leverages the strengths of the expert model but removes the limitations of the amateur model; in practice, it might for example lean toward generations that are less repetitive, which the amateur model tends to favor. Note, however, that there are probability ranges where the scaled down amateur would be preferred over the expert—e.g., when $p^{t}_{\text{EXP}}=0$, then our formula would still assign the scaled probability of the amateur to $h_{t}$. With $\gamma=1$, ContrastScore would be indifferent between the expert and the amateur but indiscriminately favor disagreement between them, an undesirable behavior similar to the division score above.